OpenStudy (anonymous):

what do i do to find x..? 5(x+6)=8 , then i got 5x+30=8x but now what

OpenStudy (phi):

5(x+6)=8 , then i got 5x+30=8x the left side is good, but the right side looks wrong. it should be 5x+30=8 (not 8x) add -30 to both sides: write + -30 on each side of the = sign and simplify

OpenStudy (anonymous):

my bad the problem is 5(x+6)=8X i forgot to put that in thier

OpenStudy (phi):

ok, 5x+30= 8x add -5x to both sides: write + -5x on both sides of the = sign and simplify

OpenStudy (anonymous):

why the negative though?

OpenStudy (phi):

try it

OpenStudy (anonymous):

30=3x x=10

OpenStudy (phi):

do you see why add -5x was helpful? It "moved" the x's all to one side

OpenStudy (anonymous):

yes but how would i know to change it to a negitive in the future?

OpenStudy (phi):

if you have 5x+30= 8x I would say "I have +5x on the left. If I add -5x to the left side, I will get +5x-5x=0, making the 5x go away. But to keep it an equation I have to add -5x to the right side (keeps things balanced)" 5x + -5x +30 = 8x -5x 30= 3x 10=x You could have add -8x to both sides, but it takes longer: 5x+ -8x +30 = 8x + -8x -3x +30 =0 -3x + 3x +30 = +3x 30= 3x 10=x

OpenStudy (anonymous):

yeah i like the first way . thanks for explaining it to me !! i thiiinkkkk i get it

OpenStudy (phi):

practice helps

OpenStudy (anonymous):

yeah it really does.

OpenStudy (anonymous):

phi , i need you help again! i have this problem . my last one . 8(4-x)+x=8+x . now . would 8+x become 9x or just 8x

OpenStudy (anonymous):

i already distributed btw . i have 32-8x+x=8+x

OpenStudy (phi):

8x+ 1x =9x (you have 8 x's and you add 1 more x to get 9 x's) 8+x is just 8+x to solve 8(4-x)+x=8+x first, distribute the 8 (multiply everything inside the parens by 8) then simplify. move the x from the right to the left side (add -x to both sides)

OpenStudy (anonymous):

okay i get that sooo i have now 32-8x+x=9x now i have to get rid of the loner x . but how

OpenStudy (phi):

no. read what I posted more carefully. 8+x is NOT 9x.

OpenStudy (anonymous):

oh . then i subtract x from both sides?

OpenStudy (phi):

yes. add -x to both sides.

OpenStudy (anonymous):

32-8x=8 ..... than i + 8 to both sides?

OpenStudy (phi):

The idea is to get x by itself on one side. I would move the numbers to the right side: add -32 to both sides

OpenStudy (anonymous):

i get x=5

OpenStudy (phi):

how?

OpenStudy (anonymous):